OpenStudy (anonymous):

f(x)=x/x+1 for (f(x+h)-f(x))/h

OpenStudy (mertsj):

\[\frac{\frac{x+h}{x+h+1}-\frac{x}{x+1}}{h}\]

OpenStudy (mertsj):

\[\frac{\frac{(x+h)(x+1)-x(x+h+1)}{(x+h+1)(x+1)}}{h}\]

OpenStudy (mertsj):

\[\frac{\frac{x^2+x+hx+h-x^2-hx-x}{(x+h+1)(x+1)}}{h}\]

OpenStudy (mertsj):

\[\frac{h}{(x+h+1)(x+1)}\times\frac{1}{h}=\frac{1}{(x+h+1)(x+1)}\]

OpenStudy (mertsj):

Are you supposed to find the limit as h approaches 0?

OpenStudy (anonymous):

no just the difference quotient

OpenStudy (anonymous):

this is the answer but I didnt understand how to do the problem thank you for explaining how you go to the answer.

OpenStudy (anonymous):

got*